OpenStudy (anonymous):

Can you show me how to solve this? 2 cos^2 (theta) - 3 cos (theta) +1 = 0

OpenStudy (anonymous):

Quadratic in cos theta (let cos theta = x)

OpenStudy (anonymous):

still need help here?

OpenStudy (anonymous):

yes please, how do you combine the cos so you can solve?

OpenStudy (anonymous):

do what @estudier said... this trigonometric equation is an equation in quadratic form... and you do know how to solve quadratic equations... using his substitution of \(\large x=cos\theta \), \[\huge 2cos^2\theta-3cos\theta+1=0\rightarrow 2x^2-3x+1=0 \] can you solve that equation in x?

OpenStudy (anonymous):

so just factor it out?

OpenStudy (anonymous):

yes... that's one way....:)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

sorry, after i factor i don't see how i can relate it to the unit circle (my answer should be \[0\le \Theta \le2\pi\]

OpenStudy (anonymous):

@dpaInc showed you already how it's done correctly, nevertheless here are the final steps in case you still have trouble, you find solutions in terms of x, therefore you have to use Back-Substitution: \[x_1=\cos \theta=1 \\x_2=\cos\theta=\frac{1}{2}\]